# CURATRON: Complete Robust Preference Data for Robust Alignment of Large Language Models

### v1

@@ 3.2 Characterization of the Adversary
Claim 1 (Upper bound on estimation error).
Under Assumption 1 it is possible that $\operatorname{dist}(\widehat{\sigma},\mathbf{P^{c}})=O(1)$.
Proof.
Assume that we are exactly given the entries of the preference matrix as opposed to sampling them. Note that in order to estimate a ranking from a given preference matrix, we still need to use a pairwise ranking procedure. Let $\widehat{\sigma}\in\mathcal{S}_{n}$ be the output of any Pairwise Ranking (PR) procedure with respect to an underlying preference matrix $\mathbf{Q}\in\mathcal{P}_{n}$. For a constant $\gamma>1$, $\widehat{\sigma}$ is said to be $\gamma$-approximate if $\operatorname{dist}(\widehat{\sigma},\mathbf{Q})\leq\gamma\min_{\sigma\in\mathcal{S}_{n}}\operatorname{dist}(\sigma,\mathbf{Q})$. Define the following distance which measures the fraction of response pairs over which two preference matrices $\{\mathbf{Q},\mathbf{R}\}\in\mathcal{P}_{n}$ disagree.
$\operatorname{dist}\left(\mathbf{Q},\mathbf{R}\right)$ $:=\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((Q_{ij}>1/2)\wedge(R_{ij}<1/2)\right)$
$+\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((Q_{ij}<1/2)\wedge(R_{ij}>1/2)\right)$
By Lemma 20 of (Rajkumar and Agarwal,, 2016), for $\mathbf{Q}\in\mathcal{P}_{n}^{ST}$ and $\mathbf{R}\in\mathcal{P}_{n}$, we have $\operatorname{dist}(\widehat{\sigma},\mathbf{Q})\leq(1+\gamma)\operatorname{dist}(\mathbf{Q},\mathbf{R})$. But note that it is possible that $\operatorname{dist}\left(\mathbf{Q},\mathbf{R}\right)=1$ as it is easy to construct by $\mathbf{R}$ that disagrees with $\mathbf{Q}$ in every entry by simply setting $\mathbf{R}=\mathbf{Q}^{\top}$. Now, we may set $\mathbf{Q}=\mathbf{P}$ and $\mathbf{R}=\mathbf{P^{c}}$ for any algorithm that uses $\mathbf{P^{c}}$ for ranking; specifically, for the adversary satisfying Assumption 1, we can see by a direct counting argument that $\operatorname{dist}\left(\mathbf{Q},\mathbf{R}\right)\leq\frac{d(2n-1-d)}{n(n-1)}$ which proves the claim. ∎
3.3 Alignment of LLMs using Direct Preference Optimization
We utilize Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) to align LLMs in our experiment to examine the efficacy of our algorithms in solving incomplete and adversarially corrupted preference data problems. DPO is a widely used method for PL, due to its simplicity, stability, performance, and computational efficiency. Essentially, DPO is a loss function that can be expressed as a mathematical equivalent to Reinforcement Learning with Human Feedback (RLHF). Its purpose is to optimize preferences by classifying human preference data without resorting to reward modeling or reinforcement learning.
$\mathcal{L}_{DPO}(\theta,y_{w},y_{l})=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w})}{\pi_{\text{ref}}(y_{w})}-\beta\log\frac{\pi_{\theta}(y_{l})}{\pi_{\text{ref}}(y_{l})}\right),$
where $y_{w}$ and $y_{l}$ represent preferred and dispreferred responses respectively. $\pi_{\text{ref}}$ is the reference policy, and $\pi_{\theta}$ is the new policy.
3.4 Solution Approach
In real-world scenarios, it is common to have missing or corrupted data, especially in large-scale crowdsource data. Our proposed methods aim to take the raw corrupted data as input and output the recovered data that closely resembles the counterfactual original data. This recovered data is then input into the alignment step, as illustrated in Figure 1. To achieve this, we first transform the raw preference data into a logit preference matrix space to utilize the low-rank characteristic. Different ranking approaches, such as, pointwise, pairwise, and listwise, can be represented in this logit preference matrix space.
${{\psi}(P)=\log{\frac{P}{1-P}}},\hskip 5.0pt\text{where}\hskip 5.0ptP={\frac{e^{-w_{i}}}{e^{-w_{i}}+e^{-w_{j}}}}$
We now identify three scenarios/settings where missing and adversarially corrupted comparisons can affect the ranking of responses and tackle them in detail in subsequent sections.
  1. 1.
  2. 2.
    Partially observed and uncorrupted setting (Figure 1(b), Section 5.1) : Not all pairs of responses are compared. This scenario can happen when data quality is prioritized over quantity in the data collection process. Observing all possible comparisons can be expensive and challenging, especially when there are many LLM responses to compare. Algorithm 4 CORATRON is proposed to solve this problem.
  3. 3.
    Partially observed and adversarially corrupted setting (Figure 1(c), Section 5.2) : Both (1) not all pairs of responses are compared, and (2) some instances of the comparison results are adversarially corrupted. This scenario can happen when data quantity and quality are not met in the data collection process. This scenario will likely happen in a large crowd-sourced environment due to large-scale LLM responses and participants. Algorithm 5 CURATRON is proposed to solve this problem.
Figure 2: Three settings/scenarios where incomplete data and adversarial corruption can affect LLMs: (a) Fully observed and adversarially corrupted setting, (b) Partially observed and uncorrupted setting, and (c) Partially observed and adversarially corrupted setting.
4 Fully Observed and Adversarially Corrupted Setting
4.1 Algorithm
  1. 1.
    Robust PCA (Procedure 1): Note that Step 3 of Algorithm 3 uses a matrix low-rank plus sparse decomposition subroutine. To obtain our recovery guarantee, it is sufficient to use the robust PCA problem as a black-box method; for the precise details of this algorithm, we refer the reader to (Netrapalli et al.,, 2014). In particular, for our analysis, we use the noise-case guarantees in their paper. This is charaterized by a (strongly-polynomial) running time of $O(n^{2}r^{2}\log(1/\epsilon))$ and guarantees $\epsilon$-recovery of the component matrices under the conditions of Assumption 1 and Lemma 3.
  2. 2.
    $\gamma$-approximate pairwise ranking procedure (Subroutine 2): Step 4 of Algorithm 3 calls a constant factor approximate ranking procedure. Specifically, we use the Copeland procedure (Copeland,, 1951) which has a $5$-approximation guarantee (Coppersmith et al.,, 2006) and involves sorting the responses according to a score of response $i$ given by $\sum_{j=1}^{n}\mathbbm{1}(\overline{P}_{ij}>1/2)$.
4.2 Analysis
Algorithm 3 RORATRON: Robust Preference Data for Robust Alignment
@@ 4.2 Analysis
We begin with a useful short result followed by the statement and the proof of our main result that, with high probability, we achieve $\epsilon$–accurate ranking in polynomial time using polynomial number of samples, despite the presence of adversarial noise. Precisely, Theorem 1 and Remark 1 address Question 3; Remark 2 addresses Question 1. In this context it is noteworthy that we present the result for LR models which strictly contain the BTL model while being much more general (Rajkumar and Agarwal,, 2016); upon proving this result, we specialize it to the classic BTL model as well (Corollary 1).
@@
Proof.
Both follow by using the definition of the logit function that $\psi(a)=\log(a/(1-a))$ and using the property that $\log(ab)=\log(a)+\log(b)$. ∎
@@
Let $\mathbf{P}\in\mathcal{P}_{n}^{LR(\psi,r)}$ be the true preference matrix according to which the pairwise comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ is generated for all responses pairs $(i,j)$ such that $k\in[K]$. Let $\mathbf{\widehat{P}}$ be the empirical preference matrix computed using $\aleph$. Let $\mathbf{S}\in[0,1]^{n\times n}$ be the adversarial matrix that additively corrupts $\mathbf{\widehat{P}}$. Let $\psi$ be $L$-Lipschitz in $[\frac{P_{\min}}{2},1-\frac{P_{\min}}{2}]$ and $\psi(\mathbf{P})$ be $\mu$-incoherent. Let each pair be compared independently $K\geq 16384\mu^{2}(1+\gamma)L^{2}n^{2}\log^{2}(n)/\epsilon\Delta^{2}$ times where $\Delta=\min_{i\neq j}\left\lvert\psi(P_{ij})-\psi(1/2)\right\rvert$. Then, with probability at least $1-1/n^{3}$, Algorithm 3 returns an estimated permutation $\widehat{\sigma}$ such that $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
@@
Note that the adversarial sparse noise we consider subsumes the setting when comparison results for certain pairs are missing as in (Rajkumar and Agarwal,, 2016) and hence directly applies in that situation. Moreover, since the support and magnitude of the corrupted entries of the preference matrix are unknown, the problem considered in this paper is harder; consequently, our sample complexity is $O(n^{2})$ as opposed to $O(n\operatorname{poly}\log n)$ in their work.
Proof.
Let $\widetilde{P}_{ij}$ be the empirical probability estimate of $P_{ij}$. Note that we compute $\widetilde{P}_{ij}=\frac{1}{K}\sum_{k=1}^{K}y_{ij}^{k}$ from the given pairwise comparison dataset, =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$. Now, $\mathbf{\widehat{P}}=\mathbf{\widetilde{P}}+\mathbf{S}$. By Lemma 1, we may write the adversarially corrupted empirical probability estimate as $\psi(\mathbf{\widehat{P}})=\psi(\mathbf{\widetilde{P}})+\mathbf{\widetilde{S}}$ where $\mathbf{\widetilde{S}}=\psi(\mathbf{\widetilde{P}}+\mathbf{S})+\psi(1-\mathbf{\widetilde{P}})$. We have $\psi(\mathbf{\widetilde{P}})=\psi(\mathbf{P})+\mathbf{\widetilde{N}}$ where $\mathbf{\widetilde{N}}=\psi(\mathbf{\widetilde{P}})-\psi(\mathbf{P})$. Now, this noise, $\mathbf{\widetilde{N}}$, is purely due to finite-sample effects which can be controlled (using concentration arguments given in the inequality $\xi_{3}$ below) by driving it down to as small a value as we want by ensuring large enough number of comparisons for each pair. Note that we input $\psi(\mathbf{\widehat{P}})=\psi(\mathbf{P})+\mathbf{\widetilde{S}}+\mathbf{\widetilde{N}}$ to Subroutine 1 and obtain $\psi(\mathbf{\overline{P}})$ as the output in Step 3 of Algorithm 3. Hence, using Theorem 2 from (Netrapalli et al.,, 2014), if $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\leq\sigma_{\min}(\psi(\mathbf{P}))/100n$, we have,
$\left\lVert\psi(\mathbf{\overline{P}})-\psi(\mathbf{P})\right\rVert_{F}\leq\epsilon^{\prime}+2\mu^{2}r\left(7\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}+\frac{8n}{r}\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\right)$
after $T\geq 10\log(3\mu^{2}r\sigma_{1}/\epsilon^{\prime})$ iterations associated with Step 1 of Subroutine  1. Next, we have, with probability at least $1-1/n^{3}$,
$\left\lVert\psi(\mathbf{\overline{P}})-\psi(\mathbf{P})\right\rVert_{F}$ $\leq\epsilon^{\prime}+2\mu^{2}r\left(7\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}+\frac{8n}{r}\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\right)$
$\stackrel{{\scriptstyle\xi_{1}}}{{\leq}}\epsilon^{\prime}+32\mu^{2}n\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\stackrel{{\scriptstyle\xi_{2}}}{{\leq}}\epsilon^{\prime}+32\mu^{2}n\tau$
$\stackrel{{\scriptstyle\xi_{3}}}{{\leq}}n\sqrt{\frac{\epsilon}{1+\gamma}}\frac{\Delta}{2}$
where $\xi_{1}$ follows by using $r\leq n$ and $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\leq\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}$, $\xi_{2}$ follows by substituting for $\mathbf{\widetilde{N}}$ from Lemma 2 with $K\geq\frac{L^{2}n^{2}\log^{2}n}{\tau^{2}}$, and $\xi_{3}$ is obtained using $\epsilon^{\prime}=n\sqrt{\frac{\epsilon}{1+\gamma}}\frac{\Delta}{4}$, $\tau=\min\left(\sigma_{\min}(\psi(\mathbf{P}))/100,\sqrt{\frac{\epsilon}{1+\gamma}}\frac{\Delta}{128\mu^{2}}\right)$. Then using similar arguments as proof of Theorem 13 in (Rajkumar and Agarwal,, 2016), we obtain our result. ∎
Lemma 2 (Concentration of sampling noise).
Under the conditions of Theorem 1, let each response pair be compared such that the number of comparisons per response pair is $K\geq\frac{L^{2}n^{2}\log(n)}{\tau^{2}}$; with probability at least $1-1/n^{3}$, $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\leq\tau$.
Proof.
Let $L$ be the Lipschitz constant of $\psi$ and set $K\geq\frac{L^{2}n^{2}\log(n)}{\tau^{2}}$. Using the inequality that $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\leq n\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}$,
$\Pr\left(\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\geq\tau\right)$ $\leq\Pr\left(\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\geq\frac{\tau}{n}\right)$
$=\Pr\left(\exists(i,j):\left\lvert\psi(\widehat{P}_{ij})-\psi(P_{ij})\right\rvert\geq\frac{\tau}{n}\right)$
$\leq\sum_{i,j}\Pr\left(\left\lvert\psi(\widehat{P}_{ij})-\psi(P_{ij})\right\rvert\geq\frac{\tau}{n}\right)$
$\leq\sum_{i,j}\Pr\left(\left\lvert\widehat{P}_{ij}-P_{ij}\right\rvert\geq\frac{\tau}{nL}\right)\leq\frac{1}{n^{3}}$
@@
The following corollary makes precise our claim that up to $O(n^{2})$ response pairs may be subject to adversarial corruption but our RORATRON algorithm still recovers a good ranking.
@@
Consider $\mathbf{P}\in\mathcal{P}_{n}^{BTL}$. Using Assumption 1, let the adversarial matrix be $\mathbf{S}\in[0,1]^{n\times n}$ satisfying $\left\lVert\mathbf{S}\right\rVert_{0}\leq n/1024\mu^{2}$ where $\mu$ is characterized as in Lemma 3. Then, with probability $1-1/n^{3}$, the output of Algorithm 3 with input $\mathbf{\widehat{P}}$ computed using =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ satisfies and $r=2$, $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
5 Partially Observed Uncorrupted and Adversarially Corrupted Settings
5.1 Partially Observed and Uncorrupted Setting
In the case when we are confident that the data are collected faithfully but we do not have the complete data to work with, we can use the ‘OptSpace’ matrix completion algorithm of (Keshavan et al.,, 2010) by itself to generate the full preference matrix from the incomplete one, as presented in Algorithm 4 CORATRON. We show in Experiment 7.3 below that we can still complete the full matrix with minimal error in such a setting with extremely missing data.
Algorithm 4 CORATRON: Complete Preference Data for Robust Alignment
0:  Comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, true rank $r$.
0:  Ranking of $n$ responses, $\hat{\sigma}\in\mathcal{S}_{n}$.
1:  Estimate entries of $\mathbf{\widehat{P}}$ for $i\leq j$ as:
2:  Set $\widehat{P}_{ij}=1-\widehat{P}_{ji}$ for all $i>j$.
4:  Using a pairwise ranking procedure after taking the inverse transform: $\widehat{\sigma}\leftarrow\text{PR}(\mathbf{R})$.
5:  return  $\widehat{\sigma}$.
5.2 Partially Observed and Adversarially Corrupted Setting
In this subsection, we consider the partially observed and adversarially corrupted comparison results setting. Both factors can be modeled in a unified manner by setting the corresponding missing entries of the preference matrix to zero (or a specific constant to account for numerical stability). We present our robust ranking algorithm for this setting in Algorithm 5 CURATRON – this essentially involves using OptSpace (Keshavan et al.,, 2010) followed by using the robust PCA algorithm of (Netrapalli et al.,, 2014) as sub-routines. We note at this point that, while (Niranjan and Rajkumar,, 2017) work considers the incomplete data case, it leverages extra information provided in the form of side information (specifically, noiseless and complete item-related features) to derive recovery guarantees. Also, their algorithm is still unable to handle the presence of pairwise comparisons corrupted in an adversarial manner, which is explored in the (Niranjan et al.,, 2017) work. We now derive the recovery guarantees as follows.
Algorithm 5 CURATRON: Complete Robust Preference Data for Robust Alignment
@@ 5.2 Partially Observed and Adversarially Corrupted Setting
Consider similar notations as in Theorem 1 but let $\mathbf{P}\in\mathcal{P}_{n}^{BTL}$. Let $\Omega\subseteq[n]\times[n]$ be a set of compared response pairs. Assume $\Omega$ is drawn uniformly from all subsets of $[n]\times[n]$ of size $\left\lvert\Omega\right\rvert$ such that $\left\lvert\Omega\right\rvert\geq C^{\prime\prime}n\log(n)$ and let the sparse noise satisfy $\left\lVert\mathbf{S}\right\rVert_{\infty}\leq\Delta_{w}\frac{\log(n)}{C_{\Delta}n}$ where $\Delta_{w}:=\min_{i,j}\left\lvert w_{i}-w_{j}\right\rvert$. Let the number of comparisons per pair be $K\geq cn^{4}/\Delta_{w}$. Then with probability at least $1-2/n^{3}$, Algorithm 5 returns a ranking that satisfies $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
@@
Proof.
From Lemma 3, we have $\psi(\mathbf{P})=\mathbf{1w}^{\top}-\mathbf{w1}^{\top}$ for the BTL model where $\psi$ is the logit function. Clearly, in this case, $\psi(\mathbf{P})$ is a real skew-symmetric matrix of rank $r=2$. Since it is skew-symmetric, its eigenvalues, which are the roots of its characteristic polynomial, are of the form ±$\pm\lambda i$ for some $\lambda\in\mathbb{R}$ and $i=\sqrt{-1}$, and hence, $\sigma_{\min}(\psi(\mathbf{P}))=\sigma_{\max}(\psi(\mathbf{P}))$, ie, the condition number of $\psi(\mathbf{P})$, $\kappa=1$. Now, we recall the spectral-lower bound from Corollary 2 of (Horne,, 1997),
$\sigma_{\min}(\psi(\mathbf{P}))\geq\frac{\left\lVert\psi(\mathbf{P})\right\rVert_{F}}{\sqrt{r(r-1)}}\geq\sqrt{\frac{n(n-1)}{2}}\Delta_{w}$ (1)
where $\Delta_{w}=\min_{i,j}\left\lvert w_{i}-w_{j}\right\rvert$.
Let $\Omega\subseteq[n]\times[n]$ be a subset of all the response pairs with comparison results among which some might be corrupted by sparse noise, ie, $\psi(\mathbf{\widehat{P}}_{\Omega})=\psi(\mathbf{P}_{\Omega})+\mathbf{\widetilde{S}}_{\Omega}+\mathbf{\widetilde{N}}_{\Omega}$. Let $\mathbf{T}:=\mathbf{\widetilde{S}}_{\Omega}+\mathbf{\widetilde{N}}_{\Omega}$. From Theorem 1.2 of (Keshavan et al.,, 2010), we have $\frac{1}{n}\left\lVert\psi(\mathbf{\widehat{P}})-\psi(\mathbf{P})\right\rVert_{F}=\frac{1}{n}\left\lVert\mathbf{T}+\mathbf{M}\right\rVert_{F}\leq C\kappa^{2}\frac{n\sqrt{r}}{\left\lvert\Omega\right\rvert}\left\lVert\mathbf{T}\right\rVert_{2}$ where $\mathbf{M}$ is the noise matrix after obtaining the completed matrix $\psi(\mathbf{\widehat{P}})$ from $\psi(\mathbf{\widehat{P}}_{\Omega})$ using OptSpace. Using triangle inequality and noting that $\left\lvert\Omega\right\rvert\geq C^{\prime\prime}n\log(n)$, the noise may be bounded as
$\left\lVert\mathbf{\widetilde{N}}_{\Omega}+\mathbf{M}\right\rVert_{\infty}$ $\leq\left\lVert\mathbf{\widetilde{N}}_{\Omega}+\mathbf{M}\right\rVert_{F}\leq\left\lVert\mathbf{T}\right\rVert_{2}\frac{\sqrt{2}Cn^{2}}{\left\lvert\Omega\right\rvert}+\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{F}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}C^{\prime}\frac{n}{\log(n)}\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}$ (2)
where $C$, $C^{\prime}$ and $C^{\prime\prime}$ are constants and $\zeta_{1}$ is obtained by using the triangle inequality that $\left\lVert\mathbf{T}\right\rVert_{2}\leq\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}+\left\lVert\mathbf{\widetilde{N}}_{\Omega}\right\rVert_{2}$, followed by setting $K\geq cn^{4}/\Delta_{w}$ for constant $c$ and finally using $\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{F}\leq\sqrt{n}\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}$. Then, combining Equations 2 and 1, we have if
$\frac{\log(n)}{C_{\Delta}n}\Delta_{w}$ $\geq\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}=\left\lVert\psi(\mathbf{\widehat{P}})-\psi(\mathbf{\widetilde{P}})\right\rVert_{2}$
$\geq\left\lVert\psi(\mathbf{\widehat{P}})-\psi(\mathbf{\widetilde{P}})\right\rVert_{\infty}\geq L\left\lVert\mathbf{\widehat{P}}-\mathbf{\widetilde{P}}\right\rVert_{\infty}\geq\left\lVert\mathbf{S}\right\rVert_{\infty}$
where $C_{\Delta}$ is a global constant and using Lemma 2, then we have the guarantee (along similar lines as that of Theorem 1 that Algorithm 5 returns an estimated permutation which satisfies $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$. ∎

### v2

@@ 4.1 Algorithm
Algorithm 3 RORATRON: Robust Preference Data for Rigorous Alignment
@@
4.2 Analysis
We begin with a useful short result followed by the statement and the Proof 13.3 of our main result that, with high probability, we achieve $\epsilon$–accurate ranking in polynomial time using polynomial number of samples, despite the presence of adversarial noise. Precisely, Theorem 1 and Remark 1 address Question 3; Remark 2 addresses Question 1. In this context, it is noteworthy that we present the result for LR models which strictly contain the BTL model while being much more general (Rajkumar and Agarwal,, 2016); upon proving this result, we specialize it to the classic BTL model as well (Corollary 1).
@@ 4.2 Analysis
Let
  1. 1.
    $\mathbf{P}\in\mathcal{P}_{n}^{LR(\psi,r)}$ be the true preference matrix according to which the pairwise comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ is generated for all responses pairs $(i,j)$ such that $k\in[K]$.
  2. 2.
    $\mathbf{\widehat{P}}$ be the empirical preference matrix computed using $\aleph$.
  3. 3.
    $\mathbf{S}\in[0,1]^{n\times n}$ be the adversarial matrix that additively corrupts $\mathbf{\widehat{P}}$, and $\mathbf{\widetilde{N}}\in\mathbf{R}^{n\times n}$ be a matrix capturing finite-sample effects.
  4. 4.
    $\psi$ be $L$-Lipschitz in $[\frac{P_{\min}}{2},1-\frac{P_{\min}}{2}]$ and $\psi(\mathbf{P})$ be $\mu$-incoherent.
  5. 5.
    Each pair be compared independently
    1. (5a)
      either $K\geq 16384\mu^{4}(1+\gamma)L^{2}\log^{2}(n)(10\log(n)+2\log 2)/\epsilon\Delta^{2}$ times when the sampling noise matrix satisfies the inequality: $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\leq 2\log(n)\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}$,
    2. (5b)
      or $K\geq 8192\mu^{4}(1+\gamma)L^{2}n^{2}(5\log(n)+\log 2)/\epsilon\Delta^{2}$ times when the sampling noise matrix satisfies the inequality: $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\leq n\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}$,
    where $\Delta=\min_{i\neq j}\left\lvert\psi(P_{ij})-\psi(1/2)\right\rvert$.
Then, with probability at least $1-1/n^{3}$, Algorithm 3 returns an estimated permutation $\widehat{\sigma}$ such that $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
@@
Note that the adversarial sparse noise we consider subsumes the setting when comparison results for certain pairs are missing as in (Rajkumar and Agarwal,, 2016) and hence directly applies in that situation. Moreover, since the support and magnitude of the corrupted entries of the preference matrix are unknown, the problem considered in this paper is harder; consequently, our sample complexity is expected to be higher, namely, $O(n^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{poly}\log n})$ (under the matrix noise model 5(5a) in Theorem 1), as opposed to $O(n\operatorname{poly}\log n)$ in their work.
@@
The following corollary makes precise our claim that up to $O(n^{2})$ response pairs may be subject to adversarial corruption, but our RORATRON algorithm still recovers a good ranking.
@@
Consider $\mathbf{P}\in\mathcal{P}_{n}^{BTL}$. Using Assumption 1, let the adversarial matrix be $\mathbf{S}\in[0,1]^{n\times n}$ satisfying $\left\lVert\mathbf{S}\right\rVert_{0}\leq n/1024\mu^{2}$ where $\mu$ is characterized as in Lemma 2. Then, with probability $1-1/n^{3}$, the output of Algorithm 3 with input $\mathbf{\widehat{P}}$ computed using =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ satisfies $r=2$ and $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
5 Partially Observed and Adversarially Corrupted Setting
In this section, we consider the partially observed and adversarially corrupted comparison results setting. Both factors can be modeled in a unified manner by setting the corresponding missing entries of the preference matrix to zero (or a specific constant to account for numerical stability). We present our robust ranking algorithm for this setting in Algorithm 4 CURATRON – this essentially involves using OptSpace (Keshavan et al.,, 2010) followed by using the RPCA algorithm of (Netrapalli et al.,, 2014) as sub-routines. We note at this point that, while (Niranjan and Rajkumar,, 2017) work considers the incomplete data case, it leverages extra information provided in the form of side information (specifically, noiseless and complete item-related features) to derive recovery guarantees. Also, their algorithm is still unable to handle the presence of pairwise comparisons corrupted in an adversarial manner, which is explored in the (Niranjan et al.,, 2017) work. We now derive the recovery guarantees as follows.
Algorithm 4 CURATRON: Complete Robust Preference Data for Rigorous Alignment
@@ 5 Partially Observed and Adversarially Corrupted Setting
Consider similar notations as in Theorem 1 but let $\mathbf{P}\in\mathcal{P}_{n}^{BTL}$. Let $\Omega\subseteq[n]\times[n]$ be a set of compared response pairs. Assume $\Omega$ is drawn uniformly from all subsets of $[n]\times[n]$ of size $\left\lvert\Omega\right\rvert$ such that $\left\lvert\Omega\right\rvert\geq C^{\prime\prime}n\log(n)$ and let the sparse noise satisfy $\left\lVert\mathbf{S}\right\rVert_{\infty}\leq\Delta_{w}\frac{\log(n)}{C_{\Delta}n}$ where $\Delta_{w}:=\min_{i,j}\left\lvert w_{i}-w_{j}\right\rvert$. Let the number of comparisons per pair be $K\geq cn^{4}/\Delta_{w}$ (under the weaker assumption on matrix noise as in 5(5b) of Theorem 1). Then with probability at least $1-2/n^{3}$, Algorithm 4 returns a ranking that satisfies $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.